# TEASMA: A Practical Methodology for Test Adequacy Assessment of Deep Neural Networks

### v3

@@ III The TEASMA Methodology
The Algorithm 1 describes the first two steps implemented in TEASMA to create an accurate regression model for predicting FDR. As illustrated in Figure 2, the first two steps of TEASMA are applied entirely on the training set of the DNN. The algorithm starts by initializing an empty archive of subset samples $A$ (line 1). Then, to be able to perform regression analysis, TEASMA is required to configure the AS computation (line 2) and identify faults in the DNN (line 3). The configuration process depends on the selected adequacy metric. In this study, it includes either generating mutants for computing MS, configuring the parameters for SC calculation, or training a VAE for IDC calculation.
@@ III-E Assessing Test Sets
In the last step, when provided with a test set, TEASMA computes its AS using the original DNN and the same adequacy metric and configuration used in previous steps (e.g., previously generated mutants). Subsequently, it predicts the test set’s FDR ($\widehat{FDR}$), including PIs based on the prediction model obtained from the previous steps. Given a certain confidence level (e.g., 95% in our experiments), such intervals specify the probabilistic range of actual FDR values for test sets achieving a specific AS value and thus provide engineers with valuable insights into the uncertainty of FDR predictions. Depending on the context, one can then decide how conservative should decision-making be by considering the lower or higher bounds of the PI. The $\widehat{FDR}$ of a test set, if deemed sufficiently accurate based on its PI, can then be used to assess its capability to detect faults and therefore whether, in a given context, it is adequate to validate the DNN.
@@ IV-B Subjects
It is important to note that our experiments were conducted using the original test sets of the subject models, as defined in their respective original sources [36, 38, 37, 39, 21], to eliminate potential biases, as we were not involved in their selection process. The subjects’ DNN models cover a wide but high accuracy range, from 77% to 96%, as detailed in Table I. As a result, scenarios where the model’s accuracy is poor, such as when a distribution shift occurs between the training and test set, are not considered in our experiments. Such scenarios reveal the need for retraining, and thus a test adequacy assessment becomes superfluous, as the goal of such assessment is to provide confidence in high-accuracy test results.
@@ IV-C AS calculation Configuration
It is important to note that not all MOs are applicable to every DNN architecture. Furthermore, the selection of MOs can affect the effectiveness of mutation analysis and thus, the results of applying TEASMA with MS. Consequently, we investigate the potential impact of MO choices in our experiments. We measure the correlation for mutants generated by each MO separately and then attempt to find subsets of operators leading to higher correlation. We compare the highest correlation achieved by subsets of MOs with that obtained with all MOs. If the correlation of the former is higher than the latter, then we only consider mutants generated by the optimal subset with the highest correlation. Detailed information regarding the applied MOs, on each subject DNN model and their respective configurations, are provided in our replication package [44].
To calculate SC, we use both the LSA and DSA metrics, consequently calculating DSC and LSC scores. When calculating DSA, first, the activation traces of training inputs for a chosen layer in the DNN are captured to establish a reference set. Then, for a given test input, the activation trace is computed at the same layer. Subsequently, the Euclidean distance between the activation trace of the test input and its nearest neighbor among the training traces is calculated, serving as the test input’s DSA score. When calculating DSA for a training input, as required in Algorithm 1, it is necessary to exclude this particular input from the training set; otherwise, the DSA value would be zero. Similarly, to accurately compute LSA for a training input, it is removed from the training set, and the probability density is recalculated based on the remaining inputs. Subsequently, LSA is calculated using the updated density estimation. This process is repeated for each training input to calculate its DSA and LSA values. In our experiments, following the approach of the original paper [6], we calculated DSA and LSA using the deepest (i.e., the last hidden) layer of the model. Additionally, we set the number of buckets for calculating LSC and DSC to 1,000, consistent with the value employed in the original paper [6]. The lower bound and upper bound for calculating the SC of both the training and test subsets are determined using the SA values observed during the training. The details of the DSC and LSC calculation configuration used for each subject are provided in our replication package [44].
@@ IV-E Data Availability
The replication package for our studies will be shared upon the paper’s acceptance [44].
@@ V Results
Before proceeding to the next steps with MS, we investigated the potential impact of MO choices when calculating MS. Ideally, MOs capable of generating mutants associated with faults are preferred in mutation analysis [17]. Therefore, it is imperative to select mutants based on which MS exhibits a stronger correlation with FDR. We calculated the correlation between FDR and MS individually for each MO as well as for different subsets of operators. Our results show that none of the MOs or MO subsets achieved a significantly higher correlation than the correlation obtained with all MOs. Consequently, we proceeded to the next step considering, for each subject, all of the applicable MOs and performed regression analysis to model the relationship between MS and FDR on the training set. Nevertheless, to provide comprehensive results, we reported the correlation for each individual MO and all MO subsets in our replication package [44].
@@
(a) $Subject\ S1$
(b) $Subject\ S2$
(c) $Subject\ S3$
(d) $Subject\ S4$
(e) $Subject\ S5$
(f) $Subject\ S6$
@@
As described in the TEASMA’s procedure in Section III, for each subject, multiple regression models are examined starting with a simple linear regression model and then using more complex non-linear models (quadratic, exponential, and regression trees) when needed. As outlined in Algorithm 1, TEASMA selects by default the regression model with the highest $R^{2}$ as the best model. In Table III, we report detailed results for the selected regression model including $R^{2}$, average RMSE, and MMRE across all $5$-folds, along with the model shape for all subjects. Further, to provide a full report, the details of all examined regression models across all subjects are reported in our replication package [44].
@@
Due to space constraints, the fitted regression curves for other metrics are provided in our replication package [44]. However, we report and discuss our observations here. When using both SC metrics, DSC, and LSC, the shape of the selected regression model is non-linear across all subjects. However, when using DSC, while the maximum FDR is consistently achieved across all subjects, the maximum DSC coverage (close to 1) is only achieved for subject S8. Moreover, the coverage of subsets with FDR close to 1 varies across subjects and is less than 0.4 for S6 and S7, less than 0.6 for S1, S2, and S5, and less than 0.7 for S3 and S4. For LSC, the maximum coverage is achieved on three subjects (S2, S4, and S6) but the coverage of subsets with FDR close to 1 varies across subjects. This indicates that achieving maximum coverage based on DSC and LSC is harder than detecting all faults. As a result, these metrics cannot provide a stopping criterion for selecting an adequate test set other than achieving the maximum coverage, which often requires selecting a large number of inputs. Consequently, this further justifies the importance of a test assessment methodology that can accurately predict a test set’s FDR. In general, the highly non-linear relationship between AS and FDR when using MS, DSC, and LSC, on most subjects, indicates that none of these metrics can be directly used to assess test sets and decide on their adequacy.
@@ V-B RQ2: Which adequacy metrics lead to better FDR predictions on test sets?
The linear regression lines between $\widehat{FDR}$ derived from MS and $ActualFDR$, where the intercept is set to zero, are shown for all subjects in Figure 5. These regression lines clearly show a strong alignment between the two. The slopes of the lines fall within the range [0.95 1.23], thus indicating no strong bias in the FDR predictions. For most subjects, the prediction models tend to slightly underestimate the actual FDR, with the exception of subject S5, where a slight overestimation is observed. Due to space constraints, the linear regression lines for the other adequacy metrics are included in our replication package [44]. However, we report and discuss our observations here. DSC exhibits the widest range of slope variations, from 0.81 for subject S5 to 1.67 for subject S3. While the prediction models tend to underestimate the actual FDR for most subjects, they overestimate it for three subjects. In contrast, IDC is the only metric where the regression slopes consistently exceed 1, ranging from 1.10 to 1.62. This suggests that for test subsets, detecting all faults is easier than achieving maximum IDC coverage.
@@ VI-B Test Adequacy Assessment
Test adequacy assessment refers to the process of evaluating the quality of a test set and its capacity to validate test outcomes [56]. In traditional software testing, coverage metrics [57] and mutation testing [56] stand out as pivotal methods for assessing test adequacy. Coverage metrics, including statement, branch, condition, and path coverage, measure the comprehensiveness of a test suite and its ability to execute different parts of the system code. Similarly, researchers have proposed various neuron coverage metrics in the context of DNN models for test assessment [8, 9, 10, 11, 20]. They aim to measure how well the test set covers the model’s internal neuron activations, ensuring that a wide range of the DNN model’s behaviors are examined during testing. However, these metrics have shown no significant correlation with the number of mispredicted inputs or detected faults in a test set [19, 58]. Consequently, in our experiments to evaluate TEASMA’s methodology with coverage metrics, we relied on SOTA surprise coverage metrics, DSC and LSC, alongside the recently introduced black-box coverage metric IDC, as described in Section II.
@@ Acknowledgements
This work was supported by a research grant from Huawei Technologies Canada Co., Ltd, as well as by the Mitacs Accelerate Program. Additional support was provided by the NSERC Discovery and Canada Research Chair programs, and the Science Foundation Ireland grant 13/RC/2094-2. The experiments conducted in this work were enabled in part by support provided by the Digital Research Alliance of Canada  [^1] https://alliancecan.ca.

### v4

@@ III The TEASMA Methodology
Algorithm 1 describes the first two steps implemented in TEASMA to create an accurate regression model for predicting FDR. As illustrated in Figure 2, the first two steps of TEASMA are applied entirely on the training set of the DNN. The algorithm starts by initializing an empty archive of subset samples $A$ (line 1). Then, to be able to perform regression analysis, TEASMA is required to configure the AS computation (line 2) and identify faults in the DNN (line 3). The configuration process depends on the selected adequacy metric. In this study, it includes either generating mutants for computing MS, configuring the parameters for SC calculation, or training a VAE for IDC calculation.
@@ III-E Assessing Test Sets
In the last step, when provided with a test set, TEASMA computes its AS using the original DNN and the same adequacy metric and configuration used in previous steps (e.g., previously generated mutants). Subsequently, it predicts the test set’s FDR ($\widehat{FDR}$), including PIs based on the prediction model obtained from the previous steps. Given a certain confidence level, such intervals specify the probabilistic range of actual FDR values for test sets achieving a specific AS value and thus provide engineers with valuable insights into the uncertainty of FDR predictions. We compute PIs for linear, quadratic, and exponential regression models, using a standard method [36]. In this method, the PIs are derived by adjusting the predicted values by a margin, which is determined by the standard error of the predictions. This adjustment accounts for the uncertainty in the estimates and includes a critical value from the t-distribution corresponding to the desired confidence level (95% in our experiments). The PIs for quadratic and exponential are constructed similarly to linear regression. But in regression tree models, we compute PIs based on the bootstrapping method for non-parametric regression models [37], where the intervals are estimated non-parametrically using the percentile method, with bounds determined by the desired significance threshold (the 2.5th and 97.5th percentiles for a 95% interval in our experiments) across the aggregated predictions from multiple trees. Depending on the context, one can then decide how conservative should decision-making be by considering the lower or higher bounds of the PI. The $\widehat{FDR}$ of a test set, if deemed sufficiently accurate based on its PI, can then be used to assess its capability to detect faults and therefore whether, in a given context, it is adequate to validate the DNN.
@@ IV-B Subjects
It is important to note that our experiments were conducted using the original test sets of the subject models, as defined in their respective original sources [38, 40, 39, 41, 21], to eliminate potential biases, as we were not involved in their selection process. The subjects’ DNN models cover a wide but high accuracy range, from 77% to 96%, as detailed in Table I. As a result, scenarios where the model’s accuracy is poor, such as when a severe distribution shift occurs between the training and test set, are not considered in our experiments. Such scenarios reveal the need for retraining, and thus a test adequacy assessment becomes superfluous, as the goal of such assessment is to provide confidence in high-accuracy test results.
@@ IV-C AS calculation Configuration
It is important to note that not all MOs are applicable to every DNN architecture. Furthermore, the selection of MOs can affect the effectiveness of mutation analysis and thus, the results of applying TEASMA with MS. Consequently, we investigate the potential impact of MO choices in our experiments. We measure the correlation for mutants generated by each MO separately and then attempt to find subsets of operators leading to higher correlation. We compare the highest correlation achieved by subsets of MOs with that obtained with all MOs. If the correlation of the former is higher than the latter, then we only consider mutants generated by the optimal subset with the highest correlation.
To calculate SC, we use both the LSA and DSA metrics, consequently calculating DSC and LSC scores. When calculating DSA, first, the activation traces of training inputs for a chosen layer in the DNN are captured to establish a reference set. Then, for a given test input, the activation trace is computed at the same layer. Subsequently, the Euclidean distance between the activation trace of the test input and its nearest neighbor among the training traces is calculated, serving as the test input’s DSA score. When calculating DSA for a training input, as required in Algorithm 1, it is necessary to exclude this particular input from the training set; otherwise, the DSA value would be zero. Similarly, to accurately compute LSA for a training input, it is removed from the training set, and the probability density is recalculated based on the remaining inputs. Subsequently, LSA is calculated using the updated density estimation. This process is repeated for each training input to calculate its DSA and LSA values. In our experiments, following the approach of the original paper [6], we calculated DSA and LSA using the deepest (i.e., the last hidden) layer of the model. Additionally, we set the number of buckets for calculating LSC and DSC to 1,000, consistent with the value employed in the original paper [6]. The lower bound and upper bound for calculating the SC of both the training and test subsets are determined using the SA values observed during the training.
@@ IV-E Data Availability
We have released the replication package of this paper online including the code, mutant models, detailed information on the MO subset analysis, and comprehensive results of all examined regression models for all metrics across all subjects [49].
@@ V Results
Before proceeding to the next steps with MS, we investigated the potential impact of MO choices when calculating MS. Ideally, MOs capable of generating mutants associated with faults are preferred in mutation analysis [17]. Therefore, it is imperative to select mutants based on which MS exhibits a stronger correlation with FDR. We calculated the correlation between FDR and MS individually for each MO as well as for different subsets of operators. Our results show that none of the MOs or MO subsets achieved a significantly higher correlation than the correlation obtained with all MOs. Consequently, we proceeded to the next step considering, for each subject, all of the applicable MOs and performed regression analysis to model the relationship between MS and FDR on training set.
@@
(a) $Subject\ S1$
(b) $Subject\ S2$
(c) $Subject\ S3$
(d) $Subject\ S4$
(e) $Subject\ S5$
(f) $Subject\ S6$
@@
As described in the TEASMA’s procedure in Section III, for each subject, multiple regression models are examined starting with a simple linear regression model and then using more complex non-linear models (quadratic, exponential, and regression trees) when needed. As outlined in Algorithm 1, TEASMA selects by default the regression model with the highest $R^{2}$ as the best model. In Table III, we report detailed results for the selected regression model including $R^{2}$, average RMSE, and MMRE across all $5$-folds, along with the model shape for all subjects.
@@
Our results with other investigated metrics indicate that when using both SC metrics, DSC, and LSC, the shape of the selected regression model is non-linear across all subjects. However, when using DSC, while the maximum FDR is consistently achieved across all subjects, the maximum DSC coverage (close to 1) is only achieved for subject S8. Moreover, the coverage of subsets with FDR close to 1 varies across subjects and is less than 0.4 for S6 and S7, less than 0.6 for S1, S2, and S5, and less than 0.7 for S3 and S4. For LSC, the maximum coverage is achieved on three subjects (S2, S4, and S6) but the coverage of subsets with FDR close to 1 varies across subjects. This indicates that achieving maximum coverage based on DSC and LSC is harder than detecting all faults. As a result, these metrics cannot provide a stopping criterion for selecting an adequate test set other than achieving the maximum coverage, which often requires selecting a large number of inputs. Consequently, this further justifies the importance of a test assessment methodology that can accurately predict a test set’s FDR. In general, the highly non-linear relationship between AS and FDR when using MS, DSC, and LSC, on most subjects, indicates that none of these metrics can be directly used to assess test sets and decide on their adequacy.
@@ V-B RQ2: Which adequacy metrics lead to better FDR predictions on test sets?
The linear regression lines between $\widehat{FDR}$ derived from MS and $ActualFDR$, where the intercept is set to zero, are shown for all subjects in Figure 5. These regression lines clearly show a strong alignment between the two. The slopes of the lines fall within the range [0.95 1.23], thus indicating no strong bias in the FDR predictions. For most subjects, the prediction models tend to slightly underestimate the actual FDR, with the exception of subject S5, where a slight overestimation is observed. Our linear regression results for the other adequacy metrics show that DSC exhibits the widest range of slope variations, from 0.81 for subject S5 to 1.67 for subject S3. While the prediction models tend to underestimate the actual FDR for most subjects, they overestimate it for three subjects. In contrast, IDC is the only metric where the regression slopes consistently exceed 1, ranging from 1.10 to 1.62. This suggests that for test subsets, detecting all faults is easier than achieving maximum IDC coverage.
@@ V-E Threats to Validity
V-F Practical Guidelines for Using TEASMA
In practice, when applying TEASMA, one must first build an accurate regression model, relying on post-training MOs, or one of the alternative metrics. Then, for each test set being considered, one must compute AS and then use the regression model to predict FDR. When relying on MS, previously generated mutants are used to calculate the test set’s MS. Therefore, the main cost is to build the regression model using the training set.
While TEASMA can be employed with various metrics, the choice of adequacy metric directly impacts both the accuracy of the FDR predictions and the overall computation time taken by the process, particularly for large input sets or complex DNN architectures. Our empirical results show that MS, DSC, and IDC are all reliable metrics for predicting FDR, though MS is the most accurate. However, their computational requirements vary significantly since, in our experiments, running TEASMA with MS took roughly five times longer than with DSC and about eight times longer than with IDC. Therefore, whether MS is applicable depends on the complexity of the DNN model under test and the size of its training set, as well as available computational resources and time constraints.
However, for high-criticality DNNs, such as those used in autonomous driving or healthcare, where accurate FDR predictions are crucial, MS is the recommended choice when possible. Conversely, for less critical systems or when faster test evaluations are required, IDC provides a good balance between accuracy and performance.
@@ VI-B Test Adequacy Assessment
Test adequacy assessment refers to the process of evaluating the quality of a test set and its capacity to validate test outcomes [57]. In traditional software testing, coverage metrics [58] and mutation testing [57] stand out as pivotal methods for assessing test adequacy. Coverage metrics, including statement, branch, condition, and path coverage, measure the comprehensiveness of a test suite and its ability to execute different parts of the system code. Similarly, researchers have proposed various neuron coverage metrics in the context of DNN models for test assessment [8, 9, 10, 11, 20, 59]. They aim to measure how well the test set covers the model’s internal neuron activations, ensuring that a wide range of the DNN model’s behaviors are examined during testing. However, these metrics have shown no significant correlation with the number of mispredicted inputs or detected faults in a test set [19, 60]. Consequently, in our experiments to evaluate TEASMA’s methodology with coverage metrics, we relied on SOTA surprise coverage metrics, DSC and LSC, alongside the recently introduced black-box coverage metric IDC, as described in Section II.
@@ Acknowledgements
This work was supported by a Research Grant from Huawei Technologies Canada Company, Ltd., the Science Foundation Ireland under Grant 13/RC/2094- 2, and the Canada Research Chair and Discovery Grant programs of the Natural Sciences and Engineering Research Council of Canada (NSERC).